1. Nolan Smith obtained a loan that requires him to make one payment at the end of 48 months. If the note requires a repayment at maturity of $10,000 and the proceeds were $6,200, what is the annual interest rate of this loan (compounded monthly, using tables, round to nearest percent)?

A. 12%

B. 1%

C. 6%

D. 2%

2. Rossville Corporation issued a ten-year $2,000,000 bond that had a 10 percent face interest rate that is paid semi-annually when the market interest rate was 8 percent. What are the proceeds generated by this bond issue (using tables, round to dollars)?

A. $2,271,830 proceeds

B. $2,342,016 proceeds

C. $1,359,030 proceeds

D.$2,000,000 proceeds

3.Which of the following statements about installment notes is NOT true?

A. All of the above are true.

B. With each subsequent payment, a larger portion of the payment covers the principal of the note.

C. All payments are the same over the life of the note.

D. The earlier payments will cover more interest expense than the later payments.

E. Depending on maturity dates, a portion of an installment note maybe classified as current and the rest as a long-term liability.
Answer #1

Solution 1:

Let monthly rate of interest = i

Periods = 48 months

Present value = $6,200

Future value = $10,000

$6,200 * (1+i)^48 = $10,000

(1+i)^48 = 1.612903

1+i = (1.612903)^1/48 = 1.010009

i =1%

Annual interest rate =1%*12 = 12%

Hence option A is correct.

Solution 2:

Computation of bond price
Table values are based on:
n= 20
i= 4.00%
Cash flow Table Value Amount Present Value
Par (Maturity) Value 0.45640 $2,000,000.00 $912,800
Interest (Annuity) 13.5903 $100,000.00 $1,359,030
Price of bonds $2,271,830

Hence option A is correct.

Solution 3:

The following statements are true about installment notes:

1. With each subsequent payment, a larger portion of the payment covers the principal of the note.

2. All payments are the same over the life of the note.

3. The earlier payments will cover more interest expense than the later payments.

4. Depending on maturity dates, a portion of an installment note maybe classified as current and the rest as a long-term liability.

Hence option A "All of the above are true". There is no false statement.